What is the future value of a 5-year ordinary annuity with annual payments of $ 720 , evaluated at a 10.71 percent interest rate? Enter your answer to the nearest $.01. Do not use $ or , signs in your answer. Enter your answer as a positive number.

Your credit card, upon which you make monthly payments, has a quoted annual interest rate of 13.5 . What is the effective annual interest rate? Calculate your answer to the nearest .01% Do not use the % sign in your answer. Thus .83% would be .83 rather than .83% or .0083.
Answer #1

1). FVA = Annuity x [{(1 + r)n - 1} / r]

= $720 x [{(1.1071)5 - 1} / 0.1071]

= $720 x [0.6632 / 0.1071] = $720 x 6.1920 = $4,458.22

2). EAR = [1 + (APR / 12)]12 - 1

= [1 + (0.135 / 12)]12 - 1 = 1.1437 - 1 = 0.1437, or 14.37